OpenStudy (emilyhoot123):

Terry is skiing down a steep hill. Terry's elevation E(t) in feet after t seconds is given by E(t) = 4500 − 80t. Write a complete sentence describing Terry's starting elevation and how it is changing over time.

OpenStudy (emilyhoot123):

does it ascend or descend

OpenStudy (kovan):

terry's elevation is descending as more time goes on

OpenStudy (saadsword):

plug in values for t

OpenStudy (saadsword):

then see if it ascends or descends

OpenStudy (emilyhoot123):

i don't know how to do that

OpenStudy (emilyhoot123):

and i need to insert the feet and feet per second

OpenStudy (saadsword):

4500- 80t t is the time in seconds. time always increases and can never be negative 4500- 80(1) is the elevation after 1 second 4500- 80(2) is the elevation after 2 seconds ans son

OpenStudy (saadsword):

**and so on**

OpenStudy (emilyhoot123):

so what do i put or need to do

OpenStudy (saadsword):

you have to Write a complete sentence describing Terry's starting elevation and how it is changing over time. Terry's starting elevation is 4500- 80(0)=4500, since 0 seconds passed since she started skiing down the hill. Terry's elevation is decreasing or descending as time goes on. this can be seen after 5 seconds since terry started going skiing down the hill: 4500- 80(5)= 4100. 4100(elevation after 5 seconds) < 4500(starting elevation) and so her elevation is decreasing or descending

OpenStudy (saadsword):

hope that helps

OpenStudy (emilyhoot123):

still confused.

OpenStudy (emilyhoot123):

ddid you see the picture i attached.

OpenStudy (saadsword):

oh

OpenStudy (saadsword):

To find the rate or feet per second use Distance = Rate*time or d=rt

OpenStudy (saadsword):

She starts at a feet of 4500

OpenStudy (emilyhoot123):

so would i put that in on the first thing

OpenStudy (saadsword):

yes

OpenStudy (saadsword):

she goes down 80 feet per second

OpenStudy (emilyhoot123):

so that means it descends ?

OpenStudy (saadsword):

yep

OpenStudy (emilyhoot123):

Find the coordinates of the midpoint of the line segment that joins the two given points. (x, y) = (−53, 19) and (x, y) = (19, −38)

OpenStudy (saadsword):

Midpiont formula : ((x1+x2)/2, (y1+y2)/2)

OpenStudy (saadsword):

plug in the values